# An observational signature for extremal black holes

### v2

@@ Abstract
We consider scalar perturbations of the Reissner–Nordström family and the Kerr family. We derive a characteristic expression of the radiation field, at any given unit solid angle of future null infinity, and numerically show that its amplitude gets excited only in the extremal case. Our work, therefore, identifies an observational signature for extremal black holes. Moreover, we show that the source of the excitation is the extremal horizon instability and its magnitude is exactly equal to the conserved horizon charge.
@@ I Introduction
Perturbations of extremal black holes exhibit a horizon instability which manifests itself in the growth along the event horizon of derivatives transversal to the event horizon aretakis1 ; aretakis2 ; aretakis4 ; hj2012 . This dynamical behavior of extremal black holes originates from a characteristic conserved charge along the event horizon, known as the horizon charge $H[\psi]$, where $\psi$ is a scalar perturbation. The existence of this charge calls into question the no-hair hypothesis nohair ; nohair2 ; nohair3 for extremal black holes according to which the only externally observable classical parameters of a black hole spacetime are the mass, angular momentum and the electric charge. On the other hand, we note that it has been shown aretakis1 that all translation-invariant quantities decay in time away from the extremal horizon; hence, the conserved charge $H[\psi]$ leaves a non-decaying trace of the perturbation only along the horizon and for this reason it can be regarded as horizon hair. In other words, the charge $H[\psi]$ could be viewed as yet another classical parameter of the extremal black hole spacetime. We remark that the observational signature of the present paper relies precisely on the existence of $H[\psi]$. Our work shows that the horizon hair $H[\psi]$ is classically measurable by external observers, including far-away observers, indeed falsifying the no-hair hypothesis in the case of extremal black holes. More specifically, we note that the proposed signature is observable by a single detector at null infinity i.e. a fixed unit solid angle; as opposed to a previously published expression that requires an integral over the entire $S^{2}$ sphere aretakis5 .
@@
It is worth noting that the same notion of horizon hair $H[\psi]$ approximately applies in the case of near extremal black holes, as a transient with duration dependent on the proximity to extremality subir2 . Therefore, we expect the results of this work to similarly hold in the near extremal cases. We also remark that, under the presence of superradiance, an amplified version dejan2023 ; zimmerman1 of the horizon instability on such backgrounds due to the existence a sequence of zero-damped quasinormal modes is already known glampedakisfull ; zeni13 .
II The geometric setting
@@
  • $\tau$: the retarded time which represents the local time of any given ideal observer at $\mathcal{I}^{+}$,
  • $\vartheta=(\theta^{1},\theta^{2})$: the unit solid angle which represents the direction of any given ideal observer along $\mathcal{I}^{+}$.
@@
where $\vartheta_{0}$ is a fixed unit solid angle.
III Observational signatures in bounded regions
@@
where $g$ is the RN/Kerr metric. Observational signatures can easily be derived in bounded regions from known results. Starting first from the event horizon, recall paper4 that asymptotically as $\tau\rightarrow\infty$ we have
@@
$\psi|_{\{r=r_{0}\}}(\tau,\vartheta)\sim\begin{cases}8I^{(1)}[\psi]\cdot\frac{1}{\tau^{3}}&\text{on subextremal RN}\\ \frac{4MH}{r_{0}-M}\cdot\frac{1}{\tau^{2}}&\text{on ERN}\end{cases}$
@@
IV The observational signature $s_{\mathcal{I}^{+}}(\vartheta)$ at null infinity
Define the radiation field $\Psi$ of the scalar field $\psi$ as follows
We then define the following limits
@@ IV The observational signature $s_{\mathcal{I}^{+}}(\vartheta)$ at null infinity
Clearly, both $C_{1}(\vartheta),C_{2}(\vartheta)$ can be computed by knowing the radiation field $\Psi(\tau,\vartheta)$ along a fixed null generator of null infinity corresponding to the solid angle $\vartheta$. The following asymptotic expansion follows then for the radiation field:
$\Psi(\tau,\vartheta)=C_{1}(\vartheta)\frac{1}{\tau^{2}}+C_{2}(\vartheta)\frac{\log(\tau/M)}{\tau^{3}}+O\left(\frac{1}{\tau^{3}}\right).$
In the sub-extremal case, it has been known logasymptotics that
$C_{1}(\vartheta)=-2I^{(1)},\ \ \ C_{2}(\vartheta)=8MI^{(1)}$ (7)
where $I^{(1)}=I^{(1)}[\psi]$ is given in terms of the initial data of $\psi$ by the following expression
@@
where the above integrals are taken relative to the respective volume forms. An alternative formula yag2020 for $I^{(1)}$ allows its computation from future null infinity as follows
[…]
Note in order to compute $I^{(1)}[\psi]$ via (8) we need to integrate the radiation field $\Psi(\tau,\vartheta)$ over all solid angles $\vartheta$. Hence formula (8) does not provide an anisotropic measurement of $I^{(1)}[\psi]$. It is important however to note that for the full sub-extremal RN family we have ${C_{2}(\vartheta)}/{C_{1}(\vartheta)}=-4M$ and this ratio is independent of $\psi$ or $\vartheta$.
Switching now to the extremal RN spacetimes, previous work paper4 has shown that
$C_{1}(\vartheta)=4MH-2I^{(1)}$ (9)
One of the numerical results of the present paper is the following expression for $C_{2}(\vartheta)$ in extremal RN:
$C_{2}(\vartheta)=-32M^{2}H+8MI^{(1)}$ (10)
for all angles $\vartheta$. The relations (9) and (10) imply that
@@
This motivates the following definition for extremal RN
@@
The signature $s_{\mathcal{I}^{+}}(\vartheta)$ allows us to differentiate extremality from sub-extremality and in fact to compute (in the extremal case) the horizon hair $H[\psi]$ by knowing the radiation field along a fixed null generator of $\mathcal{I}^{+}$.
Remark. Our analysis relies on the fact that for the full sub-extremal RN we have $\frac{C_{2}(\vartheta)}{C_{1}(\vartheta)}=-4M$. Given therefore the expression $C_{1}(\vartheta)=4MH-2I^{(1)}$ the worst-case scenario would have been that in extremal RN: $C_{2}(\vartheta)=-16M^{2}H+8MI^{(1)}$.
We can even compute $I^{(1)}$ in a similar fashion:
@@
V Numerical Approach & Results
@@
V.1 Numerical Methodology
We solve the scalar wave equation for perturbations in ERN and EK black hole backgrounds, focusing on axisymmetric modes ($m=0$). We modify the equation to work in compactified hyperboloidal coordinates $(\tau,\rho,\theta^{1},\theta^{2})$ that allow for time evolution on hypersurfaces which bring $\mathcal{I}^{+}$ to a finite radial coordinate $\rho(\mathcal{I}^{+})=S<\infty$. The relationship between these new coordinates $(\tau,\rho)$ and the spherical Boyer-Lindquist coordinates $(t,r)$ is coords
@@ V.1 Numerical Methodology
Our numerical implementation scheme entails re-writing the second order PDE in terms of two coupled first-order differential equations. We solve this system using a high-order weighted essentially non-oscillatory (WENO) finite-difference scheme with explicit Shu-Osher time-stepping. In the RN case, we solve a (1+1)D system using the standard decomposition in spherical harmonics, while in the Kerr case we solve (2+1)D system. Details may be found in our previous work weno ; subir1 ; subir2 ; subir3 ; subir4 . We choose $S=19.0$ and the location of $\mathcal{H}^{+}$ such that, $\rho(\mathcal{H}^{+})=0.95$. The initial data is a truncated Gaussian centered at $\rho=(1.0,1.1,1.2)$ with a width of $(0.16,0.22,0.32)$ and non-zero for $\rho\in[0.95,8]$. This ensures compactly supported initial data but with non-zero support on the $\mathcal{H}^{+}$ surface. The angular distribution of this initial data is spherically symmetric i.e. the $\ell=0$ monopole. The initial data’s angular distribution can be considered as “generic” i.e. includes all axisymmetric multipoles, since the higher-multipoles decay rapidly leaving the $\ell=0$ monopole to dominate at late times.
@@
V.2 Numerical Results
The numerical solution of the Teukolsky equation in compactified coordinates allows us to study the time evolution of the field at $\mathcal{I}^{+}$ with ${\theta^{1,2}}=\pi/2$. To calculate $s_{\mathcal{I}^{+}}(\vartheta)$ for subextremal RN, we first calculate the local power index (LPI) defined as $LPI\equiv\tau\dot{\psi}/\psi$ at each moment in time for $\tau/M\gtrsim 3300$. This is because under the assumptions of the previous sections for the field expansion at $\mathcal{I}^{+}$, the LPI is related to the ratio $C_{2}(\vartheta)/C_{1}(\vartheta)$ as
@@ V.2 Numerical Results
Figure 1: Time evolution of $[4M/C_{1}(\vartheta)]\times s_{\mathcal{I}^{+}}(\vartheta)$ for three sub-extremal RNs, $Q/M=[0.3,0.4,0.5]$ for any value of $\vartheta$. Note that $s_{\mathcal{I}^{+}}(\vartheta)$ asymptotes to 00 as we evolve the scalar field at $\mathcal{I}^{+}$.
@@
In Fig 1, we plot the estimated values of $s_{\mathcal{I}^{+}}(\vartheta)$ against time for three different subextremal RN cases, $Q/M=[0.3,0.4,0.5]$ at late times ($\tau/M\gtrsim 3300$). In each of these cases, it’s clear from the numerical data that $s_{\mathcal{I}^{+}}(\vartheta)$ asymptotes to 00 at late times as expected from the earlier sections. This implies that for sub-extremal BHs, irrespective of their deviation from extremality, $C_{2}(\vartheta)=-4MC_{1}(\vartheta)$ as expected.
@@
Figure 2: Best-fit line demonstrating a linear relationship between the horizon constant $H[\psi]$ and the asymptotically extracted $s_{\mathcal{I}^{+}}(\vartheta)$ for any $\vartheta$ for extremal RN. The slope of the best fit-line is $1.0019\pm 0.19$. Each data point is labelled by the initial data used for that computation i.e. the location and width of the Gaussian pulse222 The far left data point labelled by the dagger symbol uses initial data in which the field is zero while the momentum is set to the Gaussian pulse.. The Levenberg–Marquardt fitting errors in each data point are too small to be visible on the scale of the plot.
@@
Our proposal is that the quantity $s_{\mathcal{I}^{+}}(\vartheta)$ which depends on the sub-leading terms in the dynamics of the radiation field at $\mathcal{I}^{+}$ can indeed provide a signature for extremality. In order to demonstrate this, we determine $s_{\mathcal{I}^{+}}(\vartheta)$ for extremal BHs with different initial conditions and plot it against their respective constant, $H[\psi]$ that depends on initial conditions.
@@
To see the effect of sub-leading terms in the tail behavior for extremal BHs, the simulations were performed up to very late times ($t/M=8000$). Since at each step in the time evolution, we accumulate numerical error in the form of truncation of derivatives, we find that the numerical procedure employed above for sub-extremal BHs didn’t work very well for the extremal case – specifically the $LPI$ was very noisy for the extremal case. Instead, we determined $C_{1}(\vartheta)$ and $C_{2}(\vartheta)$ at $\mathcal{I}^{+}$ from the time series of the scalar field at $\mathcal{I}^{+}$ at late times ($\tau/M\gtrsim 5000$) via curve-fitting using the Levenberg–Marquardt algorithm. We plot the relationship between $s_{\mathcal{I}^{+}}(\vartheta)$ and $H[\psi]$ in Fig 2 which clearly implies a linear relationship with a slope of the best-fit line being $\approx 1$ as expected. Note that we include the origin $(0,0)$ as a constraint in our linear fitting since $s_{\mathcal{I}^{+}}(\vartheta)$ must go to zero when $H[\psi]=0$ (see the Appendix).
@@
Finally, we repeated the same analysis for Kerr black holes with axisymmetric scalar perturbations ($m=0$) focusing only on the dominant excitation mode ($l=0$). To start, in Fig 3, we plot the estimated values of $s_{\mathcal{I}^{+}}(\vartheta)$ against time for three different subextremal Kerr cases, $a/M=[0.3,0.4,0.5]$ at late times ($\tau/M\gtrsim 3300$). In each of these cases, it’s clear from the numerical data that $s_{\mathcal{I}^{+}}(\vartheta)$ asymptotes to 00 at late times.
@@
Figure 3: Time evolution of $[4M/C_{1}(\vartheta)]\times s_{\mathcal{I}^{+}}(\vartheta)$ for three sub-extremal Kerr cases, $a/M=[0.3,0.4,0.5]$ for any value of $\vartheta$. Note that $s_{\mathcal{I}^{+}}(\vartheta)$ asymptotes to 00 as we evolve the scalar field at $\mathcal{I}^{+}$.
@@
For the extremal Kerr case, we find a linear relationship between $s_{\mathcal{I}^{+}}$ (redefined below with an extra $1/4$ compared to the ERN case) vs $H[\psi]$ where the quantities are defined similar to RN black holes, i.e. in terms of $C_{1}(\vartheta)$ and $C_{2}(\vartheta)$. We present the results for EK in Fig 4. Note that $s_{\mathcal{I}^{+}}$ is redefined for Kerr with an overall $1/4$ compared to (12),
@@
$\boxed{s^{K}_{\mathcal{I}^{+}}(\vartheta)=-\frac{1}{64M^{2}}\left(C_{2}(\vartheta)+4MC_{1}(\vartheta)\right)}$ (17)

### v3

@@ Abstract
We consider scalar perturbations of the Reissner–Nordström family and the Kerr family. We derive a characteristic expression of the radiation field, at any given fixed angle of future null infinity, and numerically show that its amplitude gets excited only in the extremal case. Our work, therefore, identifies an observational signature for extremal black holes. Moreover, we show that the source of the excitation is the extremal horizon instability and its magnitude is exactly equal to the conserved horizon charge.
@@ I Introduction
Perturbations of extremal black holes exhibit a horizon instability which manifests itself in the growth along the event horizon of derivatives transversal to the event horizon aretakis1 ; aretakis2 ; aretakis4 ; hj2012 . This dynamical behavior of extremal black holes originates from a characteristic conserved charge along the event horizon, known as the horizon charge $H[\psi]$, where $\psi$ is a scalar perturbation. The existence of this charge calls into question the no-hair hypothesis nohair ; nohair2 ; nohair3 for extremal black holes according to which the only externally observable classical parameters of a black hole spacetime are the mass, angular momentum and the electric charge. On the other hand, we note that it has been shown aretakis1 that all translation-invariant quantities decay in time away from the extremal horizon; hence, the conserved charge $H[\psi]$ leaves a non-decaying trace of the perturbation only along the horizon and for this reason it can be regarded as horizon hair. In other words, the charge $H[\psi]$ could be viewed as yet another classical parameter of the extremal black hole spacetime. We remark that the observational signature of the present paper relies precisely on the existence of $H[\psi]$. This work provides numerical evidence for the measurability of the horizon hair $H[\psi]$ by external observers, including far-away observers, potentially falsifying the no-hair hypothesis in the case of extremal black holes. Note, however, that our work is in the context of an idealized isolated black hole in the framework of black hole perturbation theory i.e. a fixed background black hole spacetime. It does not apply to the case of full non-linear Einstein field equations of general relativity.
The proposed signature is observable by a single detector at null infinity i.e. a fixed angle; as opposed to a previously published expression that requires an integral over the entire $S^{2}$ sphere aretakis5 . This is significant, since it brings the in-principle observation of the hair $H[\psi]$ closer to what would be possible to do on physically realistic grounds.
The approach we take in this paper is somewhat atypical of the approach taken in the related mathematical relativity literature. We propose a conjecture which we do not attempt to prove rigorously; rather we provide compelling numerical evidence in support of it. In particular, the conjecture presented is an explicit expression for a signature $s_{\mathcal{I}^{+}}$ based on the scalar radiation field along a fixed null generator of $\mathcal{I}^{+}$. This signature is verified to be numerically zero for subextremal black holes (RN and Kerr), and shown to equal $H[\psi]$ in the extremal cases. This is the main result presented in this work. Additionally, it should be noted that while the key expressions and discussion presented is done in the context of RN black holes, the major result is also generalized to the Kerr case towards the end of the Numerical Results section.
@@
Finally, it is worth noting that the same notion of horizon hair $H[\psi]$ approximately applies in the case of near extremal black holes, as a transient with duration dependent on the proximity to extremality. Therefore, we expect the results of this work to similarly hold in the near extremal cases, especially for cases wherein the charge or spin is within $5\times 10^{-5}M$ of the extremal value subir2 . We also remark that, under the presence of superradiance, an amplified version dejan2023 ; zimmerman1 of the horizon instability on such backgrounds due to the existence a sequence of zero-damped quasinormal modes is already known glampedakisfull ; zeni13 .
II The geometric setting (RN & Kerr)
@@
  • $\tau$: the retarded time (on $\mathcal{I}^{+}$) or advanced time (on $\mathcal{H}^{+}$),
  • $\vartheta=(\theta^{1},\theta^{2})$: the angle which represents the direction of any given ideal observer along $\mathcal{I}^{+}$.
@@
where $\vartheta_{0}$ is a fixed angle.
III Summary of previous work (RN)
In this section, we briefly summarize the rigorously derived results for leading and sub-leading tail behavior for scalar fields at null infinity, $\mathcal{I}^{+}$ for sub-extremal and extremal RN black holes. In logasymptotics , the leading and sub-leading behavior of scalar fields was derived for sub-extremal RN BH’s. Specifically, this was done for four different types of initial data (ID). In this work, we will focus only on the astrophysically relevant “Type A ID” (see paper4 for definitions of ID) where the Type A ID is (i) horizon ($\mathcal{H}^{+}$) penetrating, (ii) with compact support and (iii) no support on $\mathcal{I}^{+}$.
@@ III Summary of previous work (RN)
where $g$ is the RN/Kerr metric. Following Ref. logasymptotics , we define the radiation field $\Psi$ of the scalar field $\psi$ as follows,
For Type A ID, the authors derive in logasymptotics that the asymptotic tail behavior for sub-extremal RN is given by
$\Psi(\tau,\vartheta)=C_{1}(\vartheta)\frac{1}{\tau^{2}}+C_{2}(\vartheta)\frac{\log(\tau/M)}{\tau^{3}}+O\left(\frac{1}{\tau^{3}}\right)$
They further proved that for generic non-spherically symmetric perturbations, the coefficients $C_{1}({\vartheta})$ and $C_{2}({\vartheta})$ do not depend on $\vartheta$ and only the spherically symmetric $\ell=0$ field component contributes to the value of these coefficients. Indeed the angular independence is consistent with Price’s law that states that for spherically symmetric BH space-times like RN, the leading order tail behavior at $\mathcal{I}^{+}$ for a generic perturbation with $\ell\geq 1$ is given by $1/\tau^{(\ell+2)}$ leaver .
The above results are extended to ERN in Ref. paper4 , specifically deriving late-time asymptotics for the various ID. But, for Type A ID, the authors were only able to derive the leading order behavior of $C_{1}(\vartheta)/\tau^{2}$. Additionally, they were able to prove that for the generic Type A perturbation, only the spherically symmetric component, $\ell=0$, would contribute to the coefficient $C_{1}(\vartheta)$. For Type B ID, which is compactly supported away from $\mathcal{H}^{+}$ and not supported on $\mathcal{I}^{+}$, they were able to prove that the sub-leading contribution is of the logarithmic form, $\log(\tau/M)/\tau^{3}$ and provide its coefficient explicitly. Again, the coefficient $C_{2}(\vartheta)$ is independent of $\vartheta$ on $\mathcal{I}^{+}$ and consistent with Price’s Law since $\ell\geq 1$ would contribute a term decaying as $1/\tau^{3}$ or faster.
Given the similarities between Type A/B ID, including in the vanishing of the Newman-Penrose constant $I_{0}[\psi]$we conjecture the existence of the $\log(\tau/M)/\tau^{3}$ term as the sub-leading contribution to the scalar tail behavior for both ERN and EK and further that $C_{2}(\vartheta)$, the coefficient of the $\log$ term, is dependent only on the spherically-symmetric $\ell=0$ piece of the initial perturbation motivated by Price’s Law. These taken together, prompt a proposed expression for $C_{2}(\vartheta)$ that is presented in the next section. The main contribution of this work then is to numerically verify the conjectured expression for $C_{2}(\vartheta)$ for ERN and the existence of a similar expression for EK BH’s. Finally, the existence and form of the $C_{2}(\vartheta)$ term allows us to posit an anisotropic observational signature at $\mathcal{I}^{+}$ for the extremality of BHs.
IV Observational signatures in bounded regions (RN)
Observational signatures can easily be derived in bounded regions from known results. It should be noted that unless explicitly stated otherwise, the expressions presented in this section are for RN. Starting first from the event horizon, recall paper4 that asymptotically as $\tau\rightarrow\infty$ we have
@@ IV Observational signatures in bounded regions (RN)
$\psi|_{\{r=r_{0}\}}(\tau,\vartheta)\sim\begin{cases}8I^{(1)}[\psi]\cdot\frac{1}{\tau^{3}}&\text{on subextremal RN}\\ \frac{4MH[\psi]}{r_{0}-M}\cdot\frac{1}{\tau^{2}}&\text{on ERN}\end{cases}$
@@
V The observational signature $s_{\mathcal{I}^{+}}(\vartheta)$ at null infinity (RN)
It should be noted that unless explicitly stated otherwise, the expressions presented in this section are for RN. From the radiation field $\Psi(\tau,\vartheta)$, we extract the following limits (assuming they exist)
@@ V The observational signature $s_{\mathcal{I}^{+}}(\vartheta)$ at null infinity (RN)
Clearly, both $C_{1}(\vartheta),C_{2}(\vartheta)$ can be computed by knowing the radiation field $\Psi(\tau,\vartheta)$ along a fixed null generator of null infinity corresponding to the angle $\vartheta$. In the sub-extremal case, it has been known logasymptotics that
$C_{1}(\vartheta)=-2I^{(1)}[\psi],\ \ \ C_{2}(\vartheta)=8MI^{(1)}[\psi]$ (7)
where $I^{(1)}[\psi]$ is given in terms of the initial data of $\psi$ by the following expression
@@
where the above integrals are taken relative to the respective volume forms. An alternative formula yag2020 for $I^{(1)}[\psi]$ allows its computation from future null infinity as follows
$I^{(1)}[\psi]=\frac{M}{4\pi}\int_{\mathcal{I}^{+}\cap\{\tau\geq 0\}}\Psi(\tau,\vartheta)d\vartheta d\tau.$ (8)
Note in order to compute $I^{(1)}[\psi]$ via (8) we need to integrate the radiation field $\Psi(\tau,\vartheta)$ over all angles $\vartheta$. Hence formula (8) does not provide an anisotropic measurement of $I^{(1)}[\psi]$. It is important however to note that for the full sub-extremal RN family we have ${C_{2}(\vartheta)}/{C_{1}(\vartheta)}=-4M$ and this ratio is independent of $\psi$ or $\vartheta$ logasymptotics .
Switching now to the extremal RN spacetimes, as mentioned before, previous work paper4 has shown that
$C_{1}(\vartheta)=4MH[\psi]-2I^{(1)}[\psi]$ (9)
The main conjecture that we present numerical evidence for in this work is the following expression for $C_{2}(\vartheta)$ in extremal RN:
$C_{2}(\vartheta)=-32M^{2}H[\psi]+8MI^{(1)}[\psi]$ (10)
for all angles $\vartheta$. Note that assuming $C_{2}(\vartheta)$ is a linear combination of the two numerical constants $H[\psi]$ and $I^{(1)}[\psi]$ is in part motivated by results derived for $C_{2}(\vartheta)$ for both sub-extremal RN in logasymptotics as well as sub-leading tails for Types B/C ID in paper4 . The relations (9) and (10) imply that
@@
Thus, our main conjecture motivates the following expression for a signature of extremality in the context of RN black holes
@@
The signature $s_{\mathcal{I}^{+}}(\vartheta)$ allows us to differentiate extremality from subextremality and in fact to compute (in the extremal case) the horizon hair $H[\psi]$ by knowing the radiation field along a fixed null generator of $\mathcal{I}^{+}$.
In the next section, we will present numerical evidence that supports this conjecture; in particular, we will numerically verify that $s_{\mathcal{I}^{+}}$ is zero for subextremal black holes (RN and Kerr), and numerically show that it is equal to $H[\psi]$ in the extremal cases.
Remark. Our analysis relies on the fact that for the full sub-extremal RN we have $\frac{C_{2}(\vartheta)}{C_{1}(\vartheta)}=-4M$. Given therefore the expression $C_{1}(\vartheta)=4MH[\psi]-2I^{(1)}[\psi]$ the worst-case scenario would have been that in extremal RN: $C_{2}(\vartheta)=-16M^{2}H[\psi]+8MI^{(1)}[\psi]$.
Finally, it is noteworthy that we can even compute $I^{(1)}[\psi]$ in a similar fashion:
@@
VI Numerical Approach & Results (RN & Kerr)
@@
VI.1 Numerical Methodology
We solve the scalar wave equation for perturbations in ERN and EK black hole backgrounds, focusing on axisymmetric modes ($m=0$). We modify the equation to work in compactified hyperboloidal coordinates $(\tau,\rho,\theta^{1},\theta^{2})$ that are horizon penetrating and allow for time evolution on hypersurfaces which bring $\mathcal{I}^{+}$ to a finite radial coordinate $\rho(\mathcal{I}^{+})=S<\infty$. The relationship between these new coordinates $(\tau,\rho)$ and the spherical Boyer-Lindquist coordinates $(t,r)$ is coords
@@ VI.1 Numerical Methodology
Our numerical implementation scheme entails re-writing the second order PDE in terms of two coupled first-order differential equations. We solve this system using a high-order weighted essentially non-oscillatory (WENO) finite-difference scheme with explicit Shu-Osher time-stepping. In the RN case, we solve a (1+1)D system using the standard decomposition in spherical harmonics, while in the Kerr case we solve (2+1)D system. Details may be found in our previous work weno ; subir1 ; subir2 ; subir3 ; subir4 . We choose $S=19.0$ and the location of $\mathcal{H}^{+}$ such that, $\rho(\mathcal{H}^{+})=0.95$. The initial data is a truncated Gaussian centered at $\rho=(1.0,1.1,1.2)$ with a width of $(0.16,0.22,0.32)$ and non-zero for $\rho\in[0.95,8]$. This ensures compactly supported initial data but with non-zero support on the $\mathcal{H}^{+}$ surface. The angular distribution of this initial data is spherically symmetric i.e. the $\ell=0$ monopole. However, the initial data’s angular distribution can be considered as “generic” i.e. includes all axisymmetric multipoles, since the higher-multipoles decay rapidly leaving the $\ell=0$ monopole to dominate at late times consistent with Price’s Law in both ERN and EK.
@@
VI.2 Numerical Results
The numerical solution of the wave equation in compactified coordinates allows us to study the time evolution of the field at $\mathcal{I}^{+}$ with ${\theta^{1,2}}=\pi/2$. It is important to note that while we note certain specific values of $\vartheta$ for numerical specificity, our results are actually independent of those values [^1] This is trivial by construction in ERN, of course, but not in EK. simply because at late times the $\ell=0$ mode dominates consistent with Price’s law. To calculate $s_{\mathcal{I}^{+}}(\vartheta)$ for subextremal RN, we first calculate the local power index (LPI) defined as $LPI\equiv\tau\dot{\psi}/\psi$ at each moment in time for $\tau/M\gtrsim 3300$. This is because under the assumptions of the previous sections for the field expansion at $\mathcal{I}^{+}$, the LPI is related to the ratio $C_{2}(\vartheta)/C_{1}(\vartheta)$ as
@@ VI.2 Numerical Results
Figure 1: Time evolution of $s_{\mathcal{I}^{+}}(\vartheta)$ for three sub-extremal RNs, $Q/M=[0.3,0.4,0.5]$ for any value of $\vartheta$. Note that $s_{\mathcal{I}^{+}}(\vartheta)$ asymptotes to 0 as we evolve the scalar field at $\mathcal{I}^{+}$.
@@
In Fig 1, we plot the estimated values of $s_{\mathcal{I}^{+}}(\vartheta)$ against time for three different subextremal RN cases, $Q/M=[0.3,0.4,0.5]$ at late times ($\tau/M\gtrsim 3300$). In each of these cases, it’s clear from the numerical data that $s_{\mathcal{I}^{+}}(\vartheta)$ asymptotes to 0 at late times as desired. This implies that for subextremal BHs, irrespective of their deviation from extremality, $C_{2}(\vartheta)=-4MC_{1}(\vartheta)$ as expected logasymptotics . Recall that in the subextremal cases $C_{2}(\vartheta)=-4MC_{1}(\vartheta)$ is known via mathematical derivation logasymptotics and here we are simply offering a numerical illustration of that. These numerical results for the subextremal cases help serve as a test of our numerical approach and thus build a level of confidence in our codes and analysis methodology.
@@
Figure 2: Best-fit line demonstrating a linear relationship between the horizon constant $H[\psi]$ and the asymptotically extracted $s_{\mathcal{I}^{+}}(\vartheta)$ for any $\vartheta$ for extremal RN. The slope of the best fit-line is $1.0\pm 0.2$. Each data point is labeled by the initial data used for that computation i.e. the location and width of the Gaussian pulse [^3] The far right data point labeled by the dagger symbol uses initial data in which the field is zero while the momentum is set to the Gaussian pulse.. The Levenberg–Marquardt fitting errors in each data point are too small to be visible on the scale of the plot.
@@
Our proposal is that the quantity $s_{\mathcal{I}^{+}}(\vartheta)$ which depends on the sub-leading terms in the dynamics of the radiation field at $\mathcal{I}^{+}$ can indeed provide a signature for extremality. In order to demonstrate this, we determine $s_{\mathcal{I}^{+}}(\vartheta)$ for ERN BHs with different initial conditions and plot it against their respective constant, $H[\psi]$ that depends on initial conditions.
@@
To see the effect of sub-leading terms in the tail behavior for ERN BHs, the simulations were performed up to very late times ($t/M=8000$). Since at each step in the time evolution, we accumulate numerical error in the form of truncation of derivatives, we find that the numerical procedure employed above for sub-extremal BHs didn’t work very well for the extremal case – specifically the $LPI$ was very noisy for the extremal case. Instead, we determined $C_{1}(\vartheta)$ and $C_{2}(\vartheta)$ at $\mathcal{I}^{+}$ from the time series of the scalar field at $\mathcal{I}^{+}$ at late times ($\tau/M\gtrsim 5000$) via curve-fitting using the Levenberg–Marquardt algorithm. We plot the relationship between $s_{\mathcal{I}^{+}}(\vartheta)$ and $H[\psi]$ in Fig 2 that suggests a linear relationship with a slope of the best-fit line being $\approx 1$ in accordance with our conjecture. Note that we include the origin $(0,0)$ as a constraint in our linear fitting since $s_{\mathcal{I}^{+}}(\vartheta)$ must go to zero when $H[\psi]=0$ (see the Appendix).
@@
Finally, we repeated the same analysis for Kerr black holes with axisymmetric scalar perturbations ($m=0$). To start, in Fig 3, we plot the estimated values of $s_{\mathcal{I}^{+}}(\vartheta)$ against time for three different subextremal Kerr cases, $a/M=[0.3,0.4,0.5]$ at late times ($\tau/M\gtrsim 3300$). In each of these cases, it’s clear from the numerical data that $s_{\mathcal{I}^{+}}(\vartheta)$ asymptotes to 0 at late times.
@@
Figure 3: Time evolution of $s_{\mathcal{I}^{+}}(\vartheta)$ for three sub-extremal Kerr cases, $a/M=[0.3,0.4,0.5]$ for any value of $\vartheta$. Note that $s^{\color[rgb]{0,0,0}Kerr}_{\mathcal{I}^{+}}(\vartheta)$ asymptotes to 0 as we evolve the scalar field at $\mathcal{I}^{+}$.
@@
For the extremal Kerr case, we present numerical evidence for a linear relationship between $s_{\mathcal{I}^{+}}$ (redefined below with an extra $1/4$ compared to the ERN case) vs $H[\psi]$ where the quantities are defined similar to RN black holes, i.e. in terms of $C_{1}(\vartheta)$ and $C_{2}(\vartheta)$. We present the results for EK in Fig 4. Note that $s_{\mathcal{I}^{+}}$ is redefined for Kerr with an overall $1/4$ compared to (12),
@@
$\boxed{s^{\color[rgb]{0,0,0}Kerr}_{\mathcal{I}^{+}}(\vartheta)=-\frac{1}{64M^{2}}\left(C_{2}(\vartheta)+4MC_{1}(\vartheta)\right)}$ (17)